A rectangular garden is to be surrounded by a walkway of constant width. The gardens dimensions are 24ft by 26 ft. The total area garden plus walkway is to be 1129ft^2. What must be the width of the walkway? round to the nearest thousandth if necessary.
The walkway must be _______ft wide.

Answer by mananth(16949)

24 by 26 feet sre the dimensions of the garden
let x be the width of walkway.
the length of plot will be 26+2x
width of plot will be 24+2x
Area of plot = (24+2x)(26+2x)=1129
624+48x+52x+4x^2=1129
4x^2+100x+624-1129=0
4x^2+100x-505=0
solve using quadratic equation
a=4, b=100, c=-505
x1= (-100+sqrt(100^2+(16*505)))/ (2*4)
=4.308
x2= (-100-sqrt(100^2+(16*505)))/ (2*4)
x2=-29.3 ignore this value.
The width will be 4.308 feet